# Feel the Difference? A Comparative Analysis of Emotional Arcs in Real and LLM-Generated CBT Sessions

### v1

@@ 6.1.5 Emotion Recovery Rate
Figure 2: Distribution of Spearman Correlation Scores between Real and Synthetic Speakers (Upper: Real vs. Synthetic Counselors; Lower: Real vs. Synthetic Clients)
In general, synthetic and real speakers showed similar ability to return to their baseline emotional states after a shift in valence (0.0325 vs. 0.0313) and arousal (0.0254 vs. 0.0261), with no significant differences observed. However, dominance showed partial differences: 60% of comparisons were significant, with a median p-value of 0.01965 and an average effect size of 0.217, suggesting that synthetic speakers may exert slightly more emotional control in this dimension.
@@
In contrast, synthetic clients exhibited stronger emotion regulation than real clients in both valence and dominance. Valence showed significant differences in all comparisons (median p-value: $1.698\times 10^{-5}$; average effect size: 0.401), while dominance showed partial significance in 50% of comparisons (median p-value: 0.04686; effect size: 0.168). No significant differences were found in arousal, suggesting similar regulation patterns in that dimension.
@@ 6.1.6 Discussion
Synthetic sessions are generally fluent and well-structured, yet the emotional expressions of speakers still differ in subtle but important ways from those observed in real therapy interactions. In the following, we discuss several important distinctions in the emotional dynamics of real and synthetic CBT dialogues based on our comparative analysis.
@@
6.2 Similarity between Real and Synthetic Speakers
In this section, we investigate how closely LLM-generated speakers—counselors and clients—mirror the emotional trajectories of real individuals by analyzing emotion arc similarity. Following the approach by Vishnubhotla et al. Vishnubhotla et al. (2024), we temporally align real and synthetic emotion arcs and compute Spearman correlation between each real and synthetic pair (i.e., real vs. synthetic counselors, real vs. synthetic clients). These correlation scores provide a measure of trajectory alignment: higher values indicate stronger similarity in emotional progression.
[…]
@@ 6.2.1 Discussion
The correlation analysis reveals that LLM-generated counselors and clients do not closely replicate the emotional trajectories of their real counterparts. Across both roles and all three VAD dimensions, the emotion arc similarity remains low, with most correlation values near zero and only a few outliers showing moderate alignment—often in the negative direction. This suggests that while LLMs may produce locally appropriate emotional expressions, they lack the nuanced emotional progression found in real therapeutic interactions. The weak alignment implies that current synthetic dialogues may fail to capture the dynamic, interpersonal emotional flow that characterizes real therapy, especially the moment-to-moment adaptation that occurs between counselor and client. Improving emotional trajectory modeling—particularly over longer conversational contexts—remains a key challenge for enhancing the realism and therapeutic value of LLM-generated dialogues.
@@ 8 Limitations
Our study is limited by the size and distribution of the real-world CBT dataset. Despite careful curation, the dataset comprises only 76 dialogues due to the difficulty of obtaining publicly available, high-quality therapy sessions. This limited sample size constrains the generalizability of our findings. Furthermore, the dataset is highly imbalanced in terms of speaker attributes: 84.2% of the clients are female, and 90.8% exhibit a positive overall attitude toward the session. These imbalances make it challenging to conduct reliable subgroup analyses (e.g., by gender or attitude) and may introduce bias into comparative evaluations. While we mitigate some of these limitations by focusing on the top three most frequent client problem types and applying stratified sampling in our comparisons with synthetic data, the representativeness of the dataset remains a constraint. Future work should explore larger and more demographically diverse datasets to validate and extend our findings.
@@ A.1 RealCBT Selection Criteria
A.2 Three Representative Examples of Emotion Arc Correlation between Real and Synthetic Clients
@@
These examples demonstrate the variability in arc similarity and emphasize that strong alignment is rare in synthetic client behavior.
@@
Number of sessions 58 Total client words count 59,580 Total counselor words count 82,618 Total word count 142,198 Average word count per session 2,461 Total Duration (min) 907.18 Avg. Duration (min) 15.64 Table 2: Descriptive statistics of RealCBT dataset (Top 3 Client Problems)
Category Subcategory Proportion (%) Count Client’s Problem Anxiety and fear 43.10 25 Self-esteem and confidence issues 32.76 19 Relationships (romantic, family, friendships) 24.14 14 Client’s Attitude Positive 93.10 54 Negative 3.45 2 Neutral 3.45 2 Client’s Gender Female 87.93 51 Male 12.07 7 Total 100.00 58 Table 3: Distribution of Top 3 Client Problems in the RealCBT dataset
Speaker Valence Arousal Dominance Mean Var. Mean Var. Mean Var. Real Dialogue 0.2334 0.1448 -0.0907 0.1161 0.0509 0.1131 Synthetic Dialogue 0.2761 0.1480 -0.0247 0.1037 0.1130 0.1154 Real Counselor 0.2302 0.1466 -0.0804 0.1148 0.0614 0.1120 Synthetic Counselor 0.3067 0.1287 -0.0240 0.0998 0.1550 0.1097 Real Client 0.2344 0.1312 -0.1006 0.1109 0.0346 0.1052 Synthetic Client 0.2339 0.1518 -0.0257 0.1012 0.0655 0.1043 Table 4: Emotion Mean and Variability in Valence, Arousal, and Dominance across RealCBT and CACTUS

### v3

@@ 6.1.5 Emotion Recovery Rate
In general, synthetic and real speakers showed similar ability to return to their baseline emotional states after a shift in valence (0.033 vs. 0.031) and arousal (0.025 vs. 0.026), with no significant differences observed. However, dominance showed partial differences: 60% of comparisons were significant, with a median p-value << 0.05 and $r$ = 0.22, suggesting that synthetic speakers may exert slightly more emotional control in this dimension.
@@
In contrast, synthetic clients exhibited stronger emotion regulation than real clients in both valence and dominance. Valence showed significant differences in all comparisons (median p-value << 0.0001 and $r$ = 0.4), while dominance showed partial significance in 50% of comparisons (median p-value << 0.05 and $r$ = 0.17). No significant differences were found in arousal, suggesting similar regulation patterns in that dimension.
6.1.6 Emotional Arc Case Study
To complement the quantitative results, we conduct a qualitative case study to illustrate how emotional dynamics manifest in real and synthetic CBT dialogues. Our analysis spans all three dimensions (valence, arousal, and dominance) for both counselors and clients, with detailed examples included in Appendix A.4 Table 1115. Here, we highlight counselor arousal (Emotion Mean) as a representative case: Table 1 presents excerpts showing how different levels of arousal appear in practice.
We selected three real and three synthetic sessions representing the highest, median, and lowest mean arousal scores among counselor utterances. Table 1 provides representative examples from each, along with interpretive commentary informed by key psychotherapy theories (e.g., emotion regulation Gross (1998), calibrated empathy Elliott et al. (2018, 2013), affective co-regulation Butler and Randall (2013), and therapeutic alliance theory Bordin (1979)).
These examples demonstrate how different levels of affective intensity appear in real versus LLM-generated therapy conversations. For instance, while both real and synthetic counselors express high arousal through supportive language, synthetic speakers tend to rely on generic praise (“That’s amazing progress”), which lacks co-regulatory intent and risks feeling generic or unearned. In contrast, real counselors often express enthusiasm through emotionally supportive reappraisals grounded in the client’s situation (e.g., “You’ve taken a huge step today…”). This reflects emotion regulation strategies that promote client resilience. At medium arousal, real counselors demonstrate calibrated empathy, engaging clients with reflective questions that validate while encouraging elaboration (e.g., “Can you walk me through what happened…?”). Synthetic counselors, however, tend to offer vague prompts (e.g., “Let’s explore that feeling”), which may miss therapeutic opportunities to deepen client reflection. In low-arousal moments, real counselors maintain interactional grounding by guiding the session forward, preserving the structure of therapeutic engagement. Synthetic counterparts often deliver flat acknowledgments like “Thanks for sharing,” which, while polite, offer little in terms of affective co-regulation or forward momentum. This qualitative analysis reinforces our quantitative findings: although LLMs can simulate affectively expressive language, their responses may lack the situational nuance, relational grounding, and interactive adaptability found in real therapeutic interactions.
@@ 6.1.7 Discussion
Synthetic sessions are generally fluent and well-structured, yet the emotional expressions of speakers still differ in subtle but important ways from those observed in real therapy interactions. In the following, we discuss several important distinctions in the emotional dynamics of real and synthetic CBT dialogues based on our analysis.
@@
To explain the observed discrepancies in emotional arcs between real and LLM-generated therapy dialogues, we identify several possible underlying factors. First, current LLMs may not be trained on emotionally grounded or therapeutically contextualized data Chung et al. (2023). General-purpose corpora lack the subtle affective trajectories and turn-by-turn emotion regulation seen in real counseling. Second, common prompting strategies often elicit exaggerated or uniform emotional expressions that fail to reflect the adaptive, client-contingent nature of real therapist responses Nudo et al. (2025). Third, LLMs typically lack mechanisms for maintaining contextual coherence over multiple turns, limiting their ability to simulate affective progression across a session Shu et al. (2025). Finally, real therapy involves ongoing co-regulation—therapists dynamically modulate their tone, intensity, and response based on client signals Butler and Randall (2013). This interactive nuance is missing in current LLM outputs, which treat utterances as isolated generations.
To mitigate these limitations, future work should explore fine-tuning LLMs on real counseling data to better capture authentic emotional arcs. Prompting strategies can also be refined to encourage adaptive rather than exaggerated affect. Moreover, integrating affect-aware generation and explicit co-regulatory modeling could allow LLMs to simulate more human-like emotional responsiveness and interactional flow. These directions offer promising steps toward enhancing the emotional fidelity of synthetic therapy data, which is critical for trustworthy use of LLMs in mental health applications.
6.2 Emotional Arc Similarity across Real and Synthetic Speakers
In this section, we investigate how closely the emotional trajectories of LLM-generated speakers align with those of real individuals.
6.2.1 Correlation Analysis
Building on the method by Vishnubhotla et al. Vishnubhotla et al. (2024), we temporally align emotion arcs and compute Spearman correlations between pairs of sessions. To capture different patterns of alignment, we include three pair types for each speaker role: Real–Real, Syn–Syn, and Real–Syn. This allows us to compare emotional arc similarity within real sessions, within synthetic sessions, and across real and synthetic sessions. Higher correlation values reflect stronger similarity in emotional progression. We report these comparisons separately for clients and counselors to highlight role-specific alignment dynamics (Table 2).
Client Real-Syn Real-Real Syn-Syn Valence Mean 0.014 0.0153 0.2151 Arousal Mean 0.020 0.0054 0.0225 Dominance Mean 0.002 -0.0047 0.0874 Counselor Real-Syn Real-Real Syn-Syn Valence Mean 0.044 0.0043 0.1365 Arousal Mean -0.011 0.0077 0.0213 Dominance Mean 0.058 0.0278 0.1418 Table 2: Correlation Analysis for Arc Alignment For clients, emotional arc similarity is weak across all pairings. Syn–Syn pairs show the highest values for valence (0.2151) and dominance (0.0874), but even these are modest, and arousal alignment remains negligible across all pair types. Both Real–Real and Real–Syn correlations are near zero or negative, indicating little consistency either within real sessions or between real and synthetic clients. These findings suggest that synthetic clients exhibit only limited internal consistency in their affective dynamics and fail to approximate the more complex, contingent emotion arcs of real clients.
For counselors, Syn–Syn correlations again yield the highest values, for valence (0.1365) and dominance (0.1418), but these are still weak. Real–Real and Real–Syn alignments are even lower, especially for valence and arousal. Notably, dominance alignment is slightly higher in Real–Syn pairs (0.058) than in Real–Real (0.0278), though the difference is small. Overall, these results indicate that LLM-generated counselors produce emotionally variable output, but without structured or relationally grounded patterns that resemble real therapeutic trajectories. Please see the three representative examples of emotion arc correlation between real and synthetic clients in Appendix A.5.
@@ 6.2.2 Discussion
Our quantitative analysis of emotional arc similarity reveals that alignment is consistently low across all session pairings, including Real–Real, Syn–Syn, and Real–Syn combinations. This outcome is particularly illuminating in several ways.
First, the low correlation between Real–Real pairs (e.g., 0.0153 for valence among clients and 0.0043 for counselors) reflects the natural heterogeneity of real-world therapy, where emotional trajectories are shaped by unique client issues, counselor styles, and the evolving goals of each session. In psychotherapy, affective expression is not scripted or uniform—rather, it is personalized, adaptive, and deeply contextual. Thus, low Real–Real correlation values serve as a baseline, not a shortcoming, and emphasize the richness and diversity of authentic human interactions.
By contrast, LLM-generated sessions show limited internal emotional consistency (e.g., Syn–Syn valence correlations of 0.2151 for clients and 0.1365 for counselors). While these scores are higher than those of Real–Real or Real–Syn pairs, they remain modest overall. This suggests that even when generating multiple synthetic sessions, current LLMs do not produce strongly consistent emotional arcs across speakers, pointing to a lack of coherent affective modeling.
Most notably, Real–Syn correlations are near zero across all emotion dimensions (e.g., valence = 0.014 for clients, 0.044 for counselors), indicating that synthetic speakers do not successfully emulate the affective trajectories of real speakers. This underscores a fundamental challenge: although LLMs can generate grammatically fluent and affectively expressive utterances, they struggle to reproduce the dynamic, situated, and role-sensitive emotional flow found in real therapy.
@@ 8 Limitations
Our study has several limitations. First, RealCBT is limited in size and exhibits subgroup imbalance, which affects generalizability. These constraints stem from significant privacy, ethical, and legal challenges in collecting and sharing real-world CBT data. Unlike domains where large-scale conversational datasets are readily available, therapy transcripts are highly sensitive, and even small-scale public datasets are rare. Second, our analysis focuses exclusively on CBT sessions and comparisons with a single synthetic dataset (CACTUS). While CBT is widely practiced and structurally well-defined, making it a suitable starting point, our findings may not extend to other therapy modalities, mental health domains, or cultural and linguistic contexts. Nevertheless, RealCBT remains one of the few publicly available, authentic CBT datasets. Its inclusion enables a rare grounded comparison with synthetic dialogues and offers valuable insights into how LLM-generated therapy diverges from real interactions. By making RealCBT available, we hope to catalyze broader collaborative efforts toward building larger, more diverse, and ethically responsible resources that can better capture the complexity of therapeutic dialogue.
Looking forward, we outline plans to extend research with RealCBT in several directions:
  • Expanding annotations to include dialogue acts, empathy markers, and turn-level emotional shifts;
  • Training and evaluating affect-aware LLMs that model emotional dynamics more faithfully;
  • Analyzing therapeutic strategies (e.g., reappraisal, validation) and their relationship to client emotional trajectories.
  • Broadening scope across therapeutic approaches (e.g., MI, psychodynamic therapy), languages, and cultural contexts, while incorporating alternative synthetic datasets to test robustness across generative sources.
Together, these steps aim to strengthen the dataset’s utility and support progress toward building LLM systems that are both emotionally grounded and clinically meaningful.
@@ A.1 RealCBT Selection Criteria
A.2 RealCBT Summary Statistics and Distributions
Table 3 provides an overview of the RealCBT dataset, while Table 4 presents its distribution across major categories. To highlight key patterns, Table 5 further summarizes the top three client issues represented in the dataset, and Table 6 shows their detailed distributions. Together, these tables offer a comprehensive view of both the overall dataset and the prevalence of the most common client problems.
A.3 UED Metrics in VAD across RealCBT and CACTUS
We adapt the Utterance Emotion Dynamics (UED) framework to compute emotion metrics from sequences of utterances. Table 7 reports the mean and variability of valence, arousal, and dominance (VAD) across RealCBT and CACTUS. More detailed results for each dimension are presented in Table 8 through Table 10. Together, these statistics highlight systematic differences between real and synthetic therapy dialogues, providing a basis for the comparative analyses in the paper.
A.4 Representative Examples of Emotional Arcs
This appendix provides representative examples of how emotional dynamics manifest in real and synthetic CBT dialogues. The examples span all three affective dimensions (valence, arousal, and dominance) for both counselors and clients. Selected cases, shown in Table 1 and Appendix Tables 1115, illustrate how real and synthetic utterances differ in affective expression across varying intensity levels.
A.5 Representative Real–Synthetic Client Emotion Arc Correlation Examples
@@
These examples illustrate the range of relational patterns that emerge when comparing emotional trajectories across real and LLM-generated client dialogues.
@@
Number of sessions 76
Total client words count 82,436
Total counselor words count 108,278
Total word count 190,714
Average word count per session 2,516
Total Duration (min) 1,224.67
Avg. Duration (min) 16.11
Table 3: Descriptive statistics of RealCBT dataset
@@
Number of sessions 58
Total client words count 59,580
Total counselor words count 82,618
Total word count 142,198
Average word count per session 2,461
Total Duration (min) 907.18
Avg. Duration (min) 15.64
Table 5: Descriptive statistics of RealCBT dataset (Top 3 Client Problems)
Category Subcategory Proportion (%) Count
Client’s Problem Anxiety and fear 43.10 25
Self-esteem and confidence issues 32.76 19
Relationships (romantic, family, friendships) 24.14 14
Client’s Attitude Positive 93.10 54
Negative 3.45 2
Neutral 3.45 2
Client’s Gender Female 87.93 51
Male 12.07 7
Total 100.00 58
Table 6: Distribution of Top 3 Client Problems in the RealCBT dataset
Speaker Valence Arousal Dominance
Mean Var. Mean Var. Mean Var.
Real Dialogue 0.2334 0.1448 -0.0907 0.1161 0.0509 0.1131
Synthetic Dialogue 0.2761 0.1480 -0.0247 0.1037 0.1130 0.1154
Real Counselor 0.2302 0.1466 -0.0804 0.1148 0.0614 0.1120
Synthetic Counselor 0.3067 0.1287 -0.0240 0.0998 0.1550 0.1097
Real Client 0.2344 0.1312 -0.1006 0.1109 0.0346 0.1052
Synthetic Client 0.2339 0.1518 -0.0257 0.1012 0.0655 0.1043
Table 7: Emotion Mean and Variability in Valence, Arousal, and Dominance across RealCBT and CACTUS
@@
Valence Level Real Counselor Snippet Synthetic Counselor Snippet
High "Okay. And if you continue to feel like you have to be perfect in a presentation, you know, how do you think that will affect your future in this career?" "That’s great to hear, Andrew. As a starting point, let’s work on identifying and challenging these thoughts when they arise."
Explanation: Connects the client’s concern to long-term goals, blending positive framing with constructive challenge. This reflects emotion regulation and supports the therapeutic alliance through future-oriented collaboration. Explanation: Provides encouragement but in a generic way, lacking calibration to the client’s immediate narrative. This reduces opportunities for co-regulation.
Medium "Let me ask you first. Logically, if you made a mistake in a presentation, would that make you feel like you were always going to make mistakes in presentations to follow?" "Understandably, strong feelings can be difficult to manage even when we recognize them as exaggerated."
Explanation: Uses logical probing to reduce overgeneralization, showing calibrated empathy by validating but gently challenging the client. Builds the therapeutic alliance through guided reasoning. Explanation: Normalizes emotions broadly but remains detached from the client’s specific context, limiting opportunities for emotion regulation or alliance building.
Low "Where do you feel like this fear comes from? Is there an instance that occurred that caused you to be afraid or." "Sometimes our mind can amplify fears beyond what’s likely to happen. What might be some reasons or evidence that contradict this fear?"
Explanation: Invites exploration of root causes, a low-valence but engaged stance that sustains interaction. Explanation: Provides cognitive reframing with minimal emotional tone. While constructive, it lacks the relational grounding of calibrated questioning.
Table 11: Illustrative Counselor Utterances Comparison at Varying Valence Levels
Dominance Level Real Counselor Snippet Synthetic Counselor Snippet
High "You’ve been wrestling for three years, and it is a sport you love participating in. Does it make sense that you said you were going to quit just a few matches into the season?" "I’m glad to hear you are open to it. Let’s continue to explore and challenge these beliefs together."
Explanation: Directly confronts the client’s inconsistency, showing high directive control and authority in guiding the dialogue. Explanation: Encourages collaboration but avoids strong challenge, reflecting lower dominance and shared control.
Medium "Okay, so losing is not fun. No one likes to lose. But it doesn’t mean you’re a bad wrestler." "That sounds really challenging. It must be hard to feel like you’re constantly judged. Can you recall any recent instances where you felt this way?"
Explanation: Reframes the issue and provides reassurance while still shaping the client’s perspective — moderate counselor dominance. Explanation: Invites the client to elaborate on experiences, giving more control to the client and reducing directive influence.
Low "Let me stop you right there. What if you do lose your next match? What is the worst thing that could happen?" "Let’s break it down step by step: When you noticed people staring, what specifically did you observe, and what did you think led to those thoughts and feelings?"
Explanation: Uses hypothetical exploration to reduce anxiety; less directive than a firm challenge, reflecting lower dominance. Explanation: Breaks the concern into steps, but with minimal counselor control, shifting responsibility to the client – lowest dominance.
Table 12: Illustrative Counselor Utterances Comparison at Varying Dominance Levels
Valence Level Real Client Snippet Synthetic Client Snippet
High "Not in a loving way, I guess, but that you should do your best, but also get the best. And I agree with that. I think that’s the right thing to think." "I can do that. I’ll keep a record of those moments and I’m looking forward to our next session."
Explanation: Expresses agreement through reflective reasoning, positive but context-grounded. Explanation: Optimistic and cooperative, but formulaic and less tied to personal context.
Medium "No, because I said their opinion. It was just in that time and place that I was stupid." "Hi, I’m doing okay, I guess. Just really anxious about some things."
Explanation: Self-critical and context-specific, showing moderate negative valence. Explanation: Mild disclosure of anxiety, less intense and more generic in tone.
Low "Not really? I’ve always disliked it, and I’ve had bad experiences that probably made my fear worse, but it’s. It’s always been somewhat of a fear of mine since I was a child." "I just had this fear that I would end up feeling deprived or unsatisfied, like I’d somehow starve because I couldn’t satisfy my cravings."
Explanation: Reflects long-term fear, personally grounded in past experiences, conveying sustained low valence. Explanation: Negative affect framed in exaggerated hypotheticals, less grounded in lived history.
Table 13: Illustrative Client Utterances Comparison at Varying Valence Levels
Arousal Level Real Client Snippet Synthetic Client Snippet
High "I really want to go out with y’all, but I just feel like I need to work on my. On my studies, so. I’m sorry." "I had invested a substantial amount of money, so the stress was pretty high right from the start."
Explanation: Expresses conflict and relational tension, revealing socially grounded high arousal. Explanation: Conveys stress directly but in a more factual, less relational manner.
Medium "It’s like peer pressure, sort of. But I can’t say no. I feel like I have to say yes just. Just because they are my friends. I do care about them and I don’t want to lose them." "Maybe I could see it as a learning experience rather than a total failure. It doesn’t cancel out my other achievements, right?"
Explanation: Shows interpersonal anxiety and fear of rejection, reflecting moderate arousal. Explanation: Rational reframing of failure, affectively calmer and less pressured than the real example.
Low "So I just kind of want to build up my self esteem and my confidence and accepting myself." "Sure. I’ve been feeling really unsuccessful and criticized because I’ve been taking losses in the stock market. It’s been affecting my mood and confidence a lot."
Explanation: Calm, constructive goal-setting, consistent with low arousal. Explanation: Expresses discouragement with subdued tone.
Table 14: Illustrative Client Utterances Comparison at Varying Arousal Levels
Dominance Level Real Client Snippet Synthetic Client Snippet High "My parents don’t really have the money to pay for my college, so I need to get good grades and do well with wrestling in order to have a better chance at getting a scholarship." "Thanks for having me. Well, I’ve been struggling with feeling judged because of my tattoos. It’s really been eating at me, especially in social situations." Explanation: Shows strong agency and responsibility, reflecting high dominance. Explanation: Centers on external judgment, revealing less personal control. Medium "I think about it all the time, which makes it hard for me to focus." "I think it would be helpful. I really don’t want to feel like this every time I mess up." Explanation: Rumination disrupts focus, showing partial loss of control. Explanation: Expresses a wish to change, indicating tentative but emerging agency. Low "I got so mad about losing that I wasn’t thinking straight." "Yes, just the other day I was at the grocery store, and I caught a few people staring at me. It made me so uncomfortable that I left without buying everything I needed." Explanation: Behavior dictated by anger, emotions override reasoning and control. Explanation: Actions shaped by others’ perceptions, showing externally driven loss of control. Table 15: Illustrative Client Utterances Comparison at Varying Dominance Levels